Lori wants to know which outcome (a or b) is more likely to occur: Assume there is one winning ticket in the box. Justify your answer.
a.) Ob6taining a winning ticket by drawing once from a box of 15 tickets.
b) Obtaining a winning ticket both times by drawing twice with replacement from a box of 4 tickets?

Answer by stanbon(75887)

Lori wants to know which outcome (a or b) is more likely to occur: Assume there is one winning ticket in the box. Justify your answer.
a.) Obtaining a winning ticket by drawing once from a box of 15 tickets.
P(picking the winning ticket) = 1/15
------------------
b) Obtaining a winning ticket both times by drawing twice with replacement from a box of 4 tickets?
Since you are drawing "with replacement" the probability of
drawing the winning ticket at each draw is (1/4)(1/4)=1/16
----------------
So picking the winner out of 15 has the higher probability.
Cheers,
Stan H.